OpenStudy (anonymous):

HELP WILL GIVE METAL AND FAN!!! 4. Understand the meaning and effects of arithmetic operations with integers. Please list the rules for adding, subtracting, multiplying, and dividing integers.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

addition means adding of two numbers for example 1+1=2

OpenStudy (anonymous):

thanks so is it like that for all of them @mathmath333

OpenStudy (mathmath333):

no

OpenStudy (mathmath333):

substracting means removing something from a collection for example 2-1=1 3-2=1 1-1=0

OpenStudy (mathmath333):

multiplying , we add it the number of times given for example \[4\times3=12\] \[4\times4\]=4+4+4 =8+4 =12

OpenStudy (anonymous):

thanks

OpenStudy (mathmath333):

division is splitting for example \[4\div2=2\]

OpenStudy (mathmath333):

http://www.mathsisfun.com/numbers/division.html